OpenStudy (anonymous):

State whether the given measurements determine zero, one, or two triangles. A = 61°, a = 23, b = 24

OpenStudy (raden):

first, to get the measure of angle B, use the law of sine : a sinB = b sinA sinB = b sinA/a = 24*sin61°/23 = .... ?

OpenStudy (raden):

after u got the value of sinB, take arc(sin) to get angle B and dont forget use ur calculator :)

OpenStudy (anonymous):

@RadEn calculator is broken

OpenStudy (raden):

well, in my laptop it says the value of sin61 = 0.874

OpenStudy (raden):

so, what 24 * 0.874° / 23 is = ...

OpenStudy (anonymous):

@RadEn 0.912?

OpenStudy (raden):

hehe, used calculator yea ?? :P

OpenStudy (raden):

but, that's right

OpenStudy (anonymous):

@RadEn no just used pen and paper

OpenStudy (raden):

cool, it is hard for me :)

OpenStudy (raden):

okay, then we have to fine angle B sinB = 0.192 B = arc(sin(0.192)) = 11.069 degrees = 11.07 deg (it is nearest)

OpenStudy (raden):

*find

OpenStudy (raden):

what's else ?

OpenStudy (anonymous):

@RadEn find if there is one or two triangles

OpenStudy (raden):

i think it is just one triangle, by verify the measure of angle C C = 180 - (A+B) C = 180 - (61+11.07) C = 180 - 72.07 C = 107.93 degrees only one possible for angle C

OpenStudy (raden):

do u need the length of side c ??